OpenStudy (veeveeniko):

What is the monthly payment on a $40,000 student loan at 4% interest over 15 years?

OpenStudy (cggurumanjunath):

si = ptr /100

OpenStudy (cggurumanjunath):

p=40000 t=15 r=4

OpenStudy (cggurumanjunath):

si = (40000*15*4)/100=1500

OpenStudy (amistre64):

i tend to use a formula i developed since the text formulas never really made sense to me

OpenStudy (cggurumanjunath):

a=p+i a=amount p=principal si = simple interest r= rate of interest t = time

OpenStudy (cggurumanjunath):

a= 40000+1500 =41500 dollars

OpenStudy (veeveeniko):

Oh thank you so much, I had no idea how to do this problem. You helped a lot

OpenStudy (cggurumanjunath):

rate of interest in % time in years a,p,i in dollars

OpenStudy (veeveeniko):

Wait, so the answer is 1500?

OpenStudy (cggurumanjunath):

41500

OpenStudy (amistre64):

the balance due on a loan is the compounding loan amount, minus a geometric sum of payments made. Bn = Bk^n - P(1-k^n)/(1-k) since the ending balance is 0 over n periods or 15*12 periods we have 0 = Bk^n - P(1-k^n)/(1-k) solving for P we get P = Bk^n (1-k)/(1-k^n) which can be simplified if needed to the standard text formula but i find this form easier

OpenStudy (amistre64):

k is a clean up variable for the compounding interest of (1+i/12)

OpenStudy (veeveeniko):

Wait I got confused again, how is 41,500 the answer if it's asking for the monthly payment

OpenStudy (amistre64):

filling in, i get P = 40000(1+.04/12)^(15*12) (1-(1+.04/12))/(1-(1+.04/12)^(15*12)) which the wolf or calculator easily determines

OpenStudy (amistre64):

vee, im not sure what CGG was doing, so i cant comment on their approach, but can you undertsand what i did?

OpenStudy (veeveeniko):

i think so

OpenStudy (amistre64):

taking the formula for the balance of a loan, i solved for P

OpenStudy (amistre64):

im getting 295.88 is the result of filling in the required parts

OpenStudy (veeveeniko):

I got that too

OpenStudy (amistre64):

then we agree :) yay!!

OpenStudy (veeveeniko):

I need help with another problem

OpenStudy (anonymous):

Oh that's how they did it! Thanks for posting this question. I'm curious on how to solve this problem. Now I know what to do next time! As always, helpers in OS give very clear and nice explanations! Thanks again! (: